I play Words with Friends often. It is an app which consists of a game like scrabble. I would like to calculate what my average word score is when I play a word. I have kept data from the last 30 words that I played, and my mean is 26.45 points with a standard deviation of 5.48 points. If I want to predict the population mean with 95% confidence, I need to identify what I have. What is 30?What is 5.48? What is 26.45? What equation should I use? What is my critical value? What is the confidence interval?
Answer #1

Here, 30 is the sample size

5.48 is the sample std. deviation

26.45 is the sample mean

Critical value = 2.045

Confidence interval are as:

Sample mean, xbar = 26.45
sample standard deviation, s = 5.48
sample size, n = 30
degrees of freedom, df = n - 1 = 29

Given CI level is 95%, hence α = 1 - 0.95 = 0.05
α/2 = 0.05/2 = 0.025, tc = t(α/2, df) = 2.045

CI = (26.45 - 2.045 * 5.48/sqrt(30) , 26.45 + 2.045 * 5.48/sqrt(30))
CI = (24.404 , 28.496)